Find the missing number
27   74   215
12   x     80
39  110   323

Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20060)

Find the missing number
27   74   215
12   x     80
39  110   323

We notice on the top row that
if we divide 74 by 27 we get 2.74...
and if we divide 215 by 74 we get 2.905...

We notice on the bottom row that
if we divide 110 by 39 we get 2.82...
and if we divide 323 by 110 we get 2.936...

We notice that these are all close to 3.
So the 1st and 3rd rows are close to a geometric sequence with common ratio 3.
So we check to see how much less than multiplying by 3 each next number is.

For the 1st row, 
27×3=81 we find that we must subtract 7 to get the next number, 74.
74×3=222 we find that we must subtract 7 to get the next number, 215.

For the 3rd row, 
39×3=117 we find that we must subtract 7 to get the next number, 110.
110×3=330 we find that we must subtract 7 to get the next number, 323.

So that makes us wonder if the 2nd row is like the other two.

If so then x would be 12×3=36 minus 7 or 29.

As a final check, we see what 29×3=87 minus 7 is, and hooray, it's 80.

Whoopedoo, so x=29 

Edwin

Answer by ikleyn(52835)

.
From one side, without any given context, this problem is a kind of RUBBISH.


From the other side, we see a matrix here, and the natural wish is to check if its determinant is zero.


By checking for the zero determinant, we obtain  x = 29.    ANSWER